# Magnetocalorics of singlet ground state induced moment magnets

### v1

@@ Abstract
In f-electron materials like Pr or U compounds with non-Kramers states of the f-shell the ground state may be a nonmagnetic singlet due to the action of the crystalline electric field. Nevertheless these compounds can order magnetically. They develop a ground state moment and long range order due to the spontaneous admixture of the excited state into the ground state caused by inter-site exchange. This mechanism can establish magnetic order if a control parameter exceeds a critical value defining the quantum critical point between disordered and magnetic phase. Here we investigate the magnetocaloric properties of such quantum magnets where the entropy release at low temperature is due to a competition between thermal depopulation and spontaneous order effects. We determine field and temperature dependence of specific heat for ferro- and antiferromagnetic order and also calculate the adiabatic magneto- and barocaloric cooling rates. As a model application we discuss the magnetic specific heat of the new induced ferromagnet PrI3. Furthermore we analyze the excitation spectrum of the singlet induced moment magnets in an external field with particular emphasis on field and temperature dependence of the critical soft mode. We find that the latter is fragile and exists only at specific points in the phase diagram.
@@ I Introduction
There is a class of genuine quantum magnets that cannot be described by the common semiclassical picture[1, 2] of spontaneous order of preformed moments from degenerate ground state multiplets. Most frequently these are induced moment quantum magnets with localized singlet ground states and another close by singlet or doublet, triplet state (depending on the site symmetry). The splitting is caused by the effect of the crystalline electric field (CEF) on the total angular momentum multiplets of f-electron shells, such as Pr- [3, 4, 5, 6, 7] and U- compounds [8, 9, 10, 11] with non-Kramers $f^{2}$ (J=4) configurations. The nonmagnetic ground state may nevertheless acquire an collective ordered moment by spontaneous mixing of the excited state into the ground state through the effect of the inter-site exchange. This requires the latter to exceed a critical limit as compared to the splitting energy of excited and singlet ground state. Tuning this control parameter a quantum phase transition (QPT) from paramagnetic to induced magnetic order ground state occurs. This mechanism is also possible for induced order of higher order multipolar moments like ferroquadrupolar order in 4f materials [12, 13, 14]
@@
The various possible f-electron type induced moment order has been observed in numerous Pr- and U- compounds. For a comparative investigation and examples see Ref. [15] and the references cited therein. In the present work we extend the analysis of singlet-singlet induced quantum magnetism under the application of an external field, considering ferro- (FM) as well as antiferro- (AFM) magnetic cases. We will derive the temperature- and field behavior of the induced (sublattice-) moments, and the associated H-T phase boundary as well as the critical field as function of control parameter. Particular emphasis is placed on magnetocaloric properties of the singlet-singlet induced moment magnets. We calculate field dependence of specific heat anomalies in relation to the underlying Schottky behavior compare different types of entropy release and investigate the adiabatic magnetic cooling rate. A preliminary application to the recently discovered induced moment ferromagnet PrI3 [16] is made. Furthermore we consider the analogous barocaloric cooling rate which is due to the Grüneisen parameter coupling to the CEF splitting of singlets. In particular we investigate the magnetocaloric properties in the vicinity of the quantum critical point (QCP) of the induced order and close to the critical field for the AFM. Finally we also study the dynamical properties of the singlet-singlet magnetism, deriving the temperature- and field dependence of magnetic exciton dispersion and discussing its expected (arrested) soft mode behavior in ferro- and antiferromagnetic cases and its location in the phase diagram.
@@
The ingredients of the two singlet model are defined in Sec. II. Its molecular field (MF) treatment in Sec. III leads to the selfconsistent solutions for the order parameters and renormalized singlet splitting in the ordered phase as well as critical temperature and field dependence on the control parameter. In the main Sec.  IV we derive thermodynamic potentials, specific heat and adiabatic caloric coefficients which are calculated numerically using the MF solutions. Sec. V gives a brief discussion of PrI3 in the context of the two singlet model and finally in Sec. VI we analyze in detail the field- and temperature dependent magnetic exciton dynamics and their soft mode characteristics. A discussion of numerical results and a summary are presented in Secs. VII,VIII.
@@ II The singlet-singlet CEF model in an external field
A lot of qualitative understanding of induced singlet ground state magnetism may be derived from the simple singlet-singlet CEF model. As mentioned above it has been investigated before in particular in context with Pr- metal and Pr compounds as well as U-compounds in lower than cubic, e.g. uniaxial site symmetry. In these cases the non-Kramers $4f^{2}$ or $5f^{2}$ configurations frequently lead to singlet CEF ground states. Without field its thermodynamic, quantum critical and dynamical properties have also been investigated in detail in connection with other common singlet-doublet or -triplet CEF systems [15]. The two singlet model in pseudo-spin representation may be written as
$H=\Delta\sum_{i}S_{zi}-m_{s}^{2}I_{0}\frac{1}{2}\sum_{\langle ij\rangle}S_{xi}S_{x_{j}}-m_{s}h\sum_{i}S_{xi}$ (1)
where the pseudo spin states with $S_{z}=\pm\frac{1}{2}$ correspond to the CEF ground state $|0\rangle\equiv|-\frac{1}{2}\rangle$ and $|1\rangle\equiv|+\frac{1}{2}\rangle$ at energies $-\frac{\Delta}{2}$ and $+\frac{\Delta}{2}$, respectively. Here $m_{s}=\langle 0|J_{z}|1\rangle$ is the matrix element of the physical 4f-shell total angular momentum $J_{z}$ between the two CEF states. In the $(|0\rangle,|1\rangle)$ subspace $J_{x,y}$ have not matrix elements. In the pseudo spin representation $S_{x}$ takes over the role of $J_{z}$, i.e. $J_{z}\rightarrow m_{s}S_{x}$ and likewise $S_{y}$ does not appear in the above Hamiltonian. Furthermore $h=g_{J}\mu_{B}H$ where $g_{J}$ is the Landé factor and H the strength of the external field oriented along physical (uniaxial) z-direction. The effective exchange interaction is given by $I_{e}=z|I_{0}|$ where $I_{0}$ is the nearest-neighbor (n.n.) exchange constant and z the coordination number. We use the convention $I_{0}>0$ for ferromagnetic (FM) exchange and $I_{0}<0$ for antiferromagnetic (AFM) exchange.
@@
$\hat{h}^{\lambda}_{mf}(i)$ $=$ $\Delta S_{z}(i)-m_{s}h_{e}^{\lambda}S_{x}(i)=\frac{\Delta}{2}\left(\begin{array}[]{cc}1&-\gamma^{\prime}_{\lambda}\\ -\gamma^{\prime}_{\lambda}&-1\\ \end{array}\right)$ (4)
where h is the external and $h_{e}^{\lambda}$ the effective field. Here $\lambda=A,B$ (with corresponding $(\bar{\lambda}=B,A$) denotes one of the two sublattices in the AFM case. In the FM case they are identical, i.e. $\langle S_{x}\rangle_{\lambda}\equiv\langle S_{x}\rangle$. We note that in the present model the moment is frozen along field direction by the CEF splitting $\Delta$. Therefore not only the sign of $\langle S_{x}\rangle_{A,B}$ may be opposite but their magnitude is also different for nonzero applied field. Explicitly the effective molecular field $h^{\lambda}_{e}$ or its dimensionless form $\gamma^{\prime}_{\lambda}$ is given by
@@
$\gamma^{\prime}_{\lambda}$ $=$ $\frac{m_{s}h_{e}^{\lambda}}{\Delta}=m_{s}(h^{\prime}+\sigma\gamma\langle S_{x}\rangle_{\bar{\lambda}})$ (5)
@@
Figure 1: (a) Homogeneous moment $\langle S_{x}\rangle=M_{x}^{0}$ as function of temperature for $\xi_{s}=1.2$. At zero field the ordering temperature is $T_{m}/\Delta=0.42$. At finite field the FM phase transition changes into a gradual crossover at $T_{m}$. Dashed lines show the renormalized (by order parameter and field) singlet-singlet splitting $\hat{\Delta}_{T}=\Delta_{T}/\Delta$. (b) left panel: AFM sublattice (A,B) moments, homogeneous $(M_{x}^{0})$ and staggered $(M_{x}^{s})$ moments as function of applied field $h^{\prime}=h/\Delta$ at $T/\Delta=0.01$ for a control parameter $\xi_{s}=1.2$ leading to critical field $h^{\prime}_{c}=0.77$. The induced AF phase transition persists for all fields up to $h^{\prime}_{c}$. right panel: renormalized A,B sublattice singlet splittings. The minimum $\hat{\Delta}_{B}=1$ appears when $\langle S_{x}\rangle_{B}$ changes sign (black dashed lines).
$\xi_{s}=\frac{m_{s}^{2}I_{e}}{2\Delta}$ (6)
For $\xi_{s}<1$ or $\xi_{s}>1$ the ground state will be nonmagnetic or exhibits induced magnetic (FM or AFM) order, respectively [15] with $\xi^{c}_{s}=1$ designating the quantum critical point. Alternatively we may also define $\hat{h}=m_{s}h^{\prime}=g_{\text{eff}}\mu_{B}H/\Delta$ with the effective g-factor of the two singlets given by $g_{\text{eff}}=m_{s}g_{J}$. The above Hamiltonian is readily diagonalized leading to energy levels
$\epsilon^{\pm}_{\lambda}=\pm\epsilon_{\lambda}=\pm\frac{\Delta}{2}[1+(\gamma^{\prime}_{\lambda})^{2}]^{\frac{1}{2}}$ (7)
@@
$|\epsilon^{-}_{\lambda}\rangle$ $=$ $\sin\theta_{\lambda}|1\rangle+\cos\theta_{\lambda}|0\rangle$ (8)
with the mixing angle given by $\tan 2\theta_{\lambda}=\gamma^{\prime}_{\lambda}$.
III Order parameter, magnetizations and critical temperature
@@
$\langle S_{x}\rangle_{\lambda}=\sum_{n=\pm}p^{n}_{\lambda}\langle\epsilon^{n}_{\lambda}|S_{x}|\epsilon^{n}_{\lambda}\rangle=\frac{1}{2}\tanh(\beta\epsilon_{\lambda})\sin 2\theta_{\lambda}$ (9)
@@
$=$ $[1+(m_{s}h^{\prime}+2\sigma\xi_{s}\langle S_{x}\rangle_{\bar{\lambda}})^{2}]^{\frac{1}{2}}$
Figure 2: (a) Field dependence of AFM induced moment transition temperature (phase diagram) for three different control parameters showing the increase of $T_{m}$ and $h_{c}$ with $\xi_{s}$ (full lines). For comparison the arrested soft mode frequency $\omega_{-}({\bf Q},T_{m}(h^{\prime}))$ is shown which is nonzero except at zero and critical field (dashed lines). (b) Dependence of critical temperature and critical field for the AFM transition on the control parameter above the QCP.
where $\hat{\Delta}^{\lambda}_{T}$ is the T-dependent splitting (normalized to $\Delta$) of the eigenstates on each sublattice modified due to the action of the external and molecular fields resulting from the other sublattice (for the AFM case). Naturally for FM case there is no distinction between them and the $\lambda$ index is irrelevant. In units of $(g_{\text{eff}}\mu_{B}/\text{site})$ the dimensionless homogeneous ($M^{0}_{x}$) and staggered ($M^{s}_{x}$) magnetization are then given by
$M^{0,s}_{x}=\frac{1}{2}[\langle S_{x}\rangle_{A}\pm\langle S_{x}\rangle_{B}]$ (11)
, respectively. For the FM and AFM paramagnetic $(h^{\prime}>h^{\prime}_{c})$ cases there is no sublattice dependence and we have $M_{x}^{0}=\langle S_{x}\rangle$ and $M_{x}^{s}=0$ where the homogenous magnetization at zero temperature is now obtained from Eq. (10) as solution of
$\langle S_{x}\rangle_{T=0}=M_{x}^{0}=\frac{\frac{1}{2}(m_{s}h^{\prime}+2\xi_{s}M_{x}^{0})}{[1+(m_{s}h^{\prime}+2\xi_{s}M_{x}^{0})^{2}]^{\frac{1}{2}}}$ (12)
Here the denominator is the renormalized saturated level splitting $\hat{\Delta}_{0}$. For the calculation of thermodynamic properties it will also be useful to write the selfconsistency equations in a more compact form. We will use the following identities and abbreviations:
$\gamma^{{}^{\prime}2}_{\lambda}=(\hat{\Delta}^{\lambda}_{T})^{2}-1;\;\;\;f_{s}^{\lambda}(T)=\tanh\bigl{[}(\frac{\Delta}{2T})\hat{\Delta}^{\lambda}_{T}\bigr{]}$ (13)
@@
$\langle S_{x}\rangle_{\lambda}=sign(\langle S_{x}\rangle_{\lambda})\frac{1}{2}\frac{[(\hat{\Delta}^{\lambda}_{T})^{2}-1]^{\frac{1}{2}}}{\hat{\Delta}^{\lambda}_{T}}f_{s}^{\lambda}(T)$ (14)
@@
Without external field ($h^{\prime}$= 00) the order parameter $\langle S_{x}\rangle_{\lambda}=\langle S_{x}\rangle$ is sublattice independent for FM and opposite $\langle S_{x}\rangle_{\lambda}=\lambda\langle S_{x}\rangle$ in AFM case. Then in the zero-field case Eq. (10) reduces to
$\hat{\Delta}_{T}=\xi_{s}f_{s}(T);\;\;f_{s}(T)=\tanh\bigl{[}(\frac{\Delta}{2T})\hat{\Delta}_{T}\bigr{]}$ (15)
now with simplified $\hat{\Delta}_{T}=[1+(2\xi_{s}\langle S_{x}\rangle_{T})^{2}]^{\frac{1}{2}}$ and $\langle S_{x}\rangle_{T}=\frac{1}{2\xi_{s}}[\xi_{s}^{2}f_{s}^{2}(T)-1]^{\frac{1}{2}}$ leading to a saturated splitting $\hat{\Delta}_{0}=\xi_{s}$ and moment $\langle S_{x}\rangle_{0}=\frac{1}{2\xi_{s}}(\xi_{s}^{2}-1)^{\frac{1}{2}}$ that increases with singular slope above the QCP $\xi^{c}_{s}=1$ (see Ref.15). In both cases at the transition temperature $\hat{\Delta}_{T_{m}}=1$ or $f_{s}(T_{m})=\xi_{s}^{-1}$ which leads to the explicit zero-field solution [15]:
$T_{m}(\xi_{s})$ $=$ $\frac{\Delta}{2\tanh^{-1}\bigl{(}\frac{1}{\xi_{s}}\bigr{)}}$ (16)
This means that an induced magnetic order with finite $T_{m}$ at zero field (for FM as well as AFM) appears only for $\xi_{s}>\xi_{c}=1$ which marks the QCP between paramagnetic $(\xi_{s}<\xi_{c})$ and magnetic $(\xi_{s}>\xi_{c})$ ground states. For finite field there is no transition temperature in the FM case but only a crossover behavior. In the AFM case the phase transition prevails and $T_{m}(h)$, i.e., the h-T phase diagram can be obtained from the numerical solution of Eq. (10). For FM at all $T$ and for AFM for $T>T_{m}(h^{\prime})$ $(M_{x}^{s}=0)$ the homogeneous magnetization $M_{x}^{0}(h^{\prime})$ is also obtained from this equation simply by replacing $\langle S_{x}\rangle_{\lambda},\langle S_{x}\rangle_{\bar{\lambda}}\rightarrow M_{x}^{0}$ with $\sigma=\pm$ for FM and AFM , respectively.
Figure 4: Specific heat for AFM case with $\xi_{s}=1.2$ where ordering temperature $T_{m}(h=0)$ coincides with maximum temperature $T_{max}=0.42\Delta$ of the Schottky background anomaly. For increasing field $T_{m}(h)$ decreases (Fig. 2(a)) and concomitantly the relative step size $\delta C_{V}(T_{m}(h))/C_{V}(T_{max})$ decreases (inset); they both vanish at the critical field $h_{c}(\xi_{s})$.
The order parameters, uniform and staggered magnetizations for the two cases and AFM transition temperature $T_{m}(h,\xi_{s})$ and the critical field $h_{c}(\xi_{s})$ are shown in the Figs.1,2 which will be discussed in Sec. VII.

### v2

@@ Abstract
In f-electron materials like Pr or U compounds with non-Kramers states of the f-shell the ground state may be a nonmagnetic singlet due to the action of the crystalline electric field. Nevertheless these compounds can order magnetically. They develop a ground state moment and long range order due to the spontaneous admixture of the excited state into the ground state caused by inter-site exchange. This mechanism can establish magnetic order if a control parameter exceeds a critical value defining the quantum critical point between disordered and magnetic phase. Here we investigate the magnetocaloric properties of such quantum magnets where the entropy release at low temperature is due to a competition between thermal depopulation and spontaneous order effects. We determine field and temperature dependence of specific heat for ferro- and antiferromagnetic order and also calculate the adiabatic magneto- and barocaloric cooling rates. As a model application we discuss the magnetic specific heat of the new induced ferromagnet PrIr3. Furthermore we analyze the excitation spectrum of the singlet induced moment magnets in an external field with particular emphasis on field and temperature dependence of the critical soft mode. We find that the latter is fragile and exists only at specific points in the phase diagram.
@@ I Introduction
There is a class of genuine quantum magnets that cannot be described by the common semiclassical picture[1, 2] of spontaneous order of preformed moments from degenerate ground state multiplets. Most frequently these are induced moment quantum magnets with localized singlet ground states and another close by singlet or doublet, triplet state (depending on the site symmetry). The splitting is caused by the effect of the crystalline electric field (CEF) on the total angular momentum J-multiplets of f-electron shells. Examples of such metallic f-electron materials will be discussed in later Sections. The nonmagnetic ground state may nevertheless acquire a collective ordered moment by spontaneous mixing of the excited state into the ground state through the effective RKKY inter-site exchange which is mediated by conduction electrons. For a singlet ground state the spontaneous appearance of an ordered moment requires the effective exchange to exceed a critical limit as compared to the splitting energy of ground and excited states. Increasing this control parameter beyond a critical value that defines the quantum critical point (QCP), a transition from paramagnetic to a ground state with induced magnetic order occurs. This mechanism is also possible for induced order of higher order multipolar moments like (ferro-)quadrupolar order in 4f materials.
@@
The various possible f-electron type induced moment order has been observed in numerous Pr- and U- compounds. For a comparative investigation and examples see Ref. [3] and the references cited therein. In the present work we extend the analysis of singlet-singlet induced quantum magnetism under the application of an external field, considering ferro- (FM) as well as antiferro- (AFM) magnetic cases. We will present comprehensive results firstly on the temperature- and field behavior of the induced (sublattice-) moments, the associated H-T phase boundary as well as the critical field as function of control parameter. Secondly particular emphasis is placed on a new systematic investigation of magnetocaloric properties of the singlet-singlet induced moment magnets. We calculate field dependence of specific heat jump anomalies in relation to the underlying Schottky behavior for FM as well as AFM order and demonstrate a strikingly distinct behaviour between the two cases. For AFM the relative position of the jump to the Schottky maximum position gives a measure to estimate the control parameter. Furthermore we show that the low temperature susceptibility exhibits a pronounced field dependence for control parameters close to the QCP. These predicted thermodynamic properties can be useful to characterise singlet ground state materials and as an example a preliminary application to the recently discovered induced moment ferromagnet PrIr3 [4] is made. We also compare the two different types of competing entropy release in the induced moment magnets which is not present in the semiclassical degenerate ground state magnets. As another new investigation in these systems we calculate the adiabatic magnetic cooling rate. As for the specific heat the cooling rates are very different for FM and AFM cases. In particular in the former it will be strongly enhanced for low fields and low temperature when the control parameter is slightly above the critical one. This effect should be of considerable experimental interest. Furthermore we consider the analogous barocaloric cooling rate caused by the Grüneisen parameter coupling to the CEF splitting of singlets which therefore tunes the control parameter. A large step-like anomaly at the critical pressure is predicted showing a different field dependence for FM and AFM cases.
Finally the dynamical properties of the singlet-singlet FM and AFM are studied by deriving the temperature- and field dependence of magnetic exciton dispersion. We find that the soft mode behaviour in the AFM around the field dependent transition temperature is fragile and unexpectedly becomes arrested between zero and critical field. This new prediction may be tested by inelastic neutron scattering in singlet ground state AFM.
@@
The ingredients of the two singlet model are defined in Sec. II. Its molecular field (MF) treatment in Sec. III leads to the selfconsistent solutions for the order parameters and renormalised singlet splitting in the ordered phase as well as critical temperature and field dependence on the control parameter. In the main Sec.  IV we derive thermodynamic potentials, specific heat and adiabatic caloric coefficients which are calculated numerically using the MF solutions. Sec. V gives an overview on f-electron singlet ground state materials and in particular a brief discussion of the new induced ferromagnet PrIr3 in the context of the two singlet model. Finally in Sec. VI we analyze in detail the field- and temperature dependent magnetic exciton dynamics and their soft mode characteristics. A discussion of numerical results and a summary are presented in Secs. VII,VIII.
@@ II The singlet-singlet CEF model in an external field
The localized states of 4f and 5f electron shells can approximately be characterized by a total angular momentum (J) multiplet. The action of the crystalline electric field (CEF) splits it into a sequence of CEF sub- multiplets depending on the site symmetry of the f element [5, 6]. For non-Kramers ions (integer J=4,6,8) it is possible that the lowest energy states are two singlets. This can occur generally in uniaxial or lower site symmetries [7, 3]. A list of material examples is given in Sec. V. A lot of qualitative understanding of induced singlet ground state magnetism may be derived from this simple singlet-singlet CEF model. As mentioned in the introduction it has been investigated in particular in context with (J=4) $4f^{2}$ Pr and $5f^{2}$ U intermetallic compounds in sufficiently low site symmetry. Without field the thermodynamic, quantum critical and dynamical properties of the two singlet model have also been studied in detail in connection with other common singlet-doublet or -triplet CEF systems [3]. The two singlet model in pseudo-spin representation may be written as
$H=\Delta\sum_{i}S_{zi}-m_{s}^{2}I_{0}\frac{1}{2}\sum_{\langle ij\rangle}S_{xi}S_{x_{j}}-m_{s}h\sum_{i}S_{xi},$ (1)
where the pseudo spin states with $S_{z}=\pm\frac{1}{2}$ correspond to the CEF ground state $|0\rangle\equiv|-\frac{1}{2}\rangle$ and $|1\rangle\equiv|+\frac{1}{2}\rangle$ excited state at energies $\frac{\Delta}{2}$ and $\frac{\Delta}{2}$, respectively. Here $m_{s}=\langle 0|J_{z}|1\rangle$ is the matrix element of the physical 4f-shell total angular momentum $J_{z}$ between the two CEF states. In the $(|0\rangle,|1\rangle)$ subspace $J_{x,y}$ have not matrix elements. In the pseudo spin representation $S_{x}$ takes over the role of $J_{z}$, i.e. $J_{z}\rightarrow m_{s}S_{x}$ and likewise $S_{y}$ does not appear in the above Hamiltonian. Furthermore $h=g_{J}\mu_{B}H$ where $g_{J}$ is the Landé factor and H the strength of the external field oriented along physical (uniaxial) z-direction. The effective exchange interaction is given by $I_{e}=z|I_{0}|$ where $I_{0}$ is the nearest-neighbor (n.n.) exchange constant and z the coordination number. We use the convention $I_{0}>0$ for ferromagnetic (FM) exchange and $I_{0}<0$ for antiferromagnetic (AFM) exchange.
@@
$\hat{h}^{\lambda}_{mf}(i)$ $=$ $\Delta S_{z}(i)-m_{s}h_{e}^{\lambda}S_{x}(i)$ (4)
$=$ $\frac{\Delta}{2}\left(\begin{array}[]{cc}1&-\gamma^{\prime}_{\lambda}\\ -\gamma^{\prime}_{\lambda}&-1\\ \end{array}\right),$
with h denoting the external and $h_{e}^{\lambda}$ the effective field. Here $\lambda=A,B$ (with corresponding $(\bar{\lambda}=B,A$) denotes one of the two sublattices in the AFM case. In the FM case they are identical, i.e. $\langle S_{x}\rangle_{\lambda}\equiv\langle S_{x}\rangle$. We note that in the present model the moment is frozen along field direction by the CEF splitting $\Delta$. Therefore not only the sign of $\langle S_{x}\rangle_{A,B}$ may be opposite but their magnitude is also different for nonzero applied field. Explicitly the effective molecular field $h^{\lambda}_{e}$ or its dimensionless form $\gamma^{\prime}_{\lambda}$ is defined by
@@
$\gamma^{\prime}_{\lambda}$ $=$ $\frac{m_{s}h_{e}^{\lambda}}{\Delta}=m_{s}(h^{\prime}+\sigma\gamma\langle S_{x}\rangle_{\bar{\lambda}}),$ (5)
@@
Figure 1: (a) Homogeneous moment $\langle S_{x}\rangle=M_{x}^{0}$ as function of temperature for $\xi_{s}=1.2$. At zero field the ordering temperature is $T_{m}/\Delta=0.42$. At finite field the FM phase transition changes into a gradual crossover at $T_{m}$. Dashed lines show the renormalised (by order parameter and field) singlet-singlet splitting $\hat{\Delta}_{T}=\Delta_{T}/\Delta$. (b) Left panel: AFM sublattice (A,B) moments, homogeneous $(M_{x}^{0})$ and staggered $(M_{x}^{s})$ moments as function of applied field $h^{\prime}=h/\Delta$ at $T/\Delta=0.01$ for a control parameter $\xi_{s}=1.2$ leading to critical field $h^{\prime}_{c}=0.77$. The induced AF phase transition persists for all fields up to $h^{\prime}_{c}$. Right panel: renormalised A,B sublattice singlet splittings. The minimum $\hat{\Delta}_{B}=1$ appears when $\langle S_{x}\rangle_{B}$ changes sign (black dashed lines).
$\xi_{s}=\frac{m_{s}^{2}I_{e}}{2\Delta}.$ (6)
For $\xi_{s}<1$ or $\xi_{s}>1$ the ground state will be nonmagnetic or exhibits induced magnetic (FM or AFM) order, respectively [3] with $\xi^{c}_{s}=1$ designating the quantum critical point. As a combination of various microscopic parameters of the strongly correlated f-states and their effective intersite interaction this central important quantity has to be considered as an adjustable model parameter. As we argue later it can be estimated from specific heat analysis or the dispersion characteristics of magnetic exciton modes. In some cases a tuning of the control parameter by applying external (uniaxial) pressure [8] or substitution with nonmagnetic elements [9] has been achieved.
For the dimensionless field we may alternatively define $\hat{h}=m_{s}h^{\prime}=g_{\text{eff}}\mu_{B}H/\Delta$ with the effective g-factor of the two singlets given by $g_{\text{eff}}=m_{s}g_{J}$. The above Hamiltonian $\hat{h}^{\lambda}_{mf}$ is readily diagonalised leading to energy levels
$\epsilon^{\pm}_{\lambda}=\pm\epsilon_{\lambda}=\pm\frac{\Delta}{2}[1+(\gamma^{\prime}_{\lambda})^{2}]^{\frac{1}{2}},$ (7)
@@
$|\epsilon^{-}_{\lambda}\rangle$ $=$ $\sin\theta_{\lambda}|1\rangle+\cos\theta_{\lambda}|0\rangle,$ (8)
with the mixing angle given by $\tan 2\theta_{\lambda}=\gamma^{\prime}_{\lambda}$. The constant energy shift contained in the first line of Eq. (4) is irrelevant for calculating the molecular fields but has to be included later for the internal and free energies in Sec. IV.1.
III Order parameter, magnetisations and critical temperature
@@
$\langle S_{x}\rangle_{\lambda}=\sum_{n=\pm}p^{n}_{\lambda}\langle\epsilon^{n}_{\lambda}|S_{x}|\epsilon^{n}_{\lambda}\rangle=\frac{1}{2}\tanh(\beta\epsilon_{\lambda})\sin 2\theta_{\lambda},$ (9)
@@
$=$ $[1+(m_{s}h^{\prime}+2\sigma\xi_{s}\langle S_{x}\rangle_{\bar{\lambda}})^{2}]^{\frac{1}{2}},$
Figure 2: (a) Field dependence of AFM induced moment transition temperature (phase diagram) for three different control parameters showing the increase of $T_{m}$ and $h_{c}$ with $\xi_{s}$ (full lines). For comparison the arrested soft mode frequency $\omega_{-}({\bf Q},T_{m}(h^{\prime}))$ is shown which is nonzero except at zero and critical field (dashed lines). (b) Dependence of AFM critical temperature Tm (for h=0) and critical field hc (for T=0) on the control parameter above the QCP.
where $\hat{\Delta}^{\lambda}_{T}$ is the T-dependent splitting (normalised to $\Delta$) of the eigenstates on each sublattice modified due to the action of the external and molecular fields resulting from the other sublattice (for the AFM case). Naturally for FM case there is no distinction between them and the $\lambda$ index is irrelevant. In units of $(g_{\text{eff}}\mu_{B}/\text{site})$ the dimensionless homogeneous ($M^{0}_{x}$) and staggered ($M^{s}_{x}$) magnetisation are then given by
$M^{0,s}_{x}=\frac{1}{2}[\langle S_{x}\rangle_{A}\pm\langle S_{x}\rangle_{B}],$ (11)
respectively. For the FM and AFM paramagnetic $(h^{\prime}>h^{\prime}_{c})$ cases there is no sublattice dependence and we have $M_{x}^{0}=\langle S_{x}\rangle$ and $M_{x}^{s}=0$ where the homogenous magnetisation at zero temperature is now obtained from Eq. (10) as solution of
$\langle S_{x}\rangle_{T=0}=M_{x}^{0}=\frac{\frac{1}{2}(m_{s}h^{\prime}+2\xi_{s}M_{x}^{0})}{[1+(m_{s}h^{\prime}+2\xi_{s}M_{x}^{0})^{2}]^{\frac{1}{2}}}.$ (12)
Here the denominator is the renormalised saturated level splitting $\hat{\Delta}_{0}$. For the calculation of thermodynamic properties it will also be useful to write the selfconsistency equations in a more compact form. We will use the following identities and abbreviations:
$\gamma^{{}^{\prime}2}_{\lambda}=(\hat{\Delta}^{\lambda}_{T})^{2}-1;\;\;\;f_{s}^{\lambda}(T)=\tanh\bigl{[}(\frac{\Delta}{2T})\hat{\Delta}^{\lambda}_{T}\bigr{]}.$ (13)
@@
$\langle S_{x}\rangle_{\lambda}=sign(\langle S_{x}\rangle_{\lambda})\frac{1}{2}\frac{[(\hat{\Delta}^{\lambda}_{T})^{2}-1]^{\frac{1}{2}}}{\hat{\Delta}^{\lambda}_{T}}f_{s}^{\lambda}(T),$ (14)
@@
Without external field ($h^{\prime}$= 0) the order parameter $\langle S_{x}\rangle_{\lambda}=\langle S_{x}\rangle$ is sublattice independent for FM and opposite $\langle S_{x}\rangle_{\lambda}=\lambda\langle S_{x}\rangle$ in AFM case. Then in the zero-field case Eq. (10) reduces to
$\hat{\Delta}_{T}=\xi_{s}f_{s}(T);\;\;f_{s}(T)=\tanh\bigl{[}(\frac{\Delta}{2T})\hat{\Delta}_{T}\bigr{]},$ (15)
now with simplified $\hat{\Delta}_{T}=[1+(2\xi_{s}\langle S_{x}\rangle_{T})^{2}]^{\frac{1}{2}}$ and $\langle S_{x}\rangle_{T}=\frac{1}{2\xi_{s}}[\xi_{s}^{2}f_{s}^{2}(T)-1]^{\frac{1}{2}}$ leading to a saturated splitting $\hat{\Delta}_{0}=\xi_{s}$ and moment $\langle S_{x}\rangle_{0}=\frac{1}{2\xi_{s}}(\xi_{s}^{2}-1)^{\frac{1}{2}}$ that increases with singular slope above the QCP $\xi^{c}_{s}=1$ (see Ref.3). In both cases at the transition temperature $\hat{\Delta}_{T_{m}}=1$ or $f_{s}(T_{m})=\xi_{s}^{-1}$ which leads to the explicit zero-field solution [3]:
$T_{m}(\xi_{s})$ $=$ $\frac{\Delta}{2\tanh^{-1}\bigl{(}\frac{1}{\xi_{s}}\bigr{)}}.$ (16)
This means that an induced magnetic order with finite $T_{m}$ at zero field (for FM as well as AFM) appears only for $\xi_{s}>\xi_{c}=1$ which marks the QCP between paramagnetic $(\xi_{s}<\xi_{c})$ and magnetic $(\xi_{s}>\xi_{c})$ ground states. For finite field there is no transition temperature in the FM case but only a crossover behaviour. In the AFM case the phase transition prevails and $T_{m}(h)$, i.e., the h-T phase diagram can be obtained from the numerical solution of Eq. (10). For FM at all $T$ and for AFM for $T>T_{m}(h^{\prime})$ $(M_{x}^{s}=0)$ the homogeneous magnetisation $M_{x}^{0}(h^{\prime})$ is also obtained from this equation simply by replacing $\langle S_{x}\rangle_{\lambda},\langle S_{x}\rangle_{\bar{\lambda}}\rightarrow M_{x}^{0}$ with $\sigma=\pm$ for FM and AFM , respectively.
Figure 4: (a) Specific heat for AFM case with $\xi_{s}=1.2$ where ordering temperature $T_{m}(h=0)$ coincides with maximum temperature $T_{max}=0.42\Delta$ of the Schottky background anomaly. For increasing field $T_{m}(h)$ decreases (Fig. 2(a)) and concomitantly the relative step size $\delta C_{V}(T_{m}(h))/C_{V}(T_{max})$ decreases (inset); they both vanish at the critical field $h_{c}(\xi_{s})$. (b) Specific heat for FM with $\xi_{s}=1.2$ which is identical to AFM for $h=0$. However for finite field the FM order changes rapidly into a crossover (cf. Fig. 1(a)) and the sharp jump broadens and merges with the Schottky background for larger field. This type of field dependence should be expected for the induced FM PrIr3 shown in Fig. 5 (only for zero field case). .
The order parameters, uniform and staggered magnetisations for the two cases and AFM transition temperature $T_{m}(h,\xi_{s})$ and the critical field $h_{c}(\xi_{s})$ are shown in the Figs.1,2 which will be discussed in Sec. VII.